GT Inc.’s net income before tax on its financial statements was $700,000, and its taxable income was $810,000. The $110,000 difference is the aggregate of temporary book/tax differences. GT’s tax rate is 21 percent.

  1. Compute GT’s tax expense for financial statement purposes.
  2. Compute GT’s tax payable.
  3. Compute the net increase in GT’s deferred tax assets or deferred tax liabilities (identify which) for the year.
Answer #1

GT Inc's Tax expense for financial statement purpose will be calculated on the net income before tax as reported in the financial statement.

Tax Expense = $700,000 * 21% = $147,000

GT Inc's Tax payable will be calculated on the taxable income which is calculated for tax purposes.

Tax Payable = $810,000 * 21% = $170,100

Whenever there is excess tax payment or excess tax payable over tax expense, it results in deferred tax assets.

Deferred Tax Asset = Tax Payable - Tax Expenses

Deferred Tax Asset = $170,100 - $147,000

Deferred Tax Asset = $23,100

OR,

Deferred Tax Asset = (Taxable Income - Net Income before tax reported in financial statement) * Tax Rate

(NOTE: this can be only used when the tax rate is same for both tax purpose and for net income in financial statement, if the tax rates are different then tax expense and tax payable needs to be calculated separately before calculation of deferred tax aasets or deferred tax liabilities)

Deferred Tax Asset = ($810,000 - $700,000)* 21% = $23,100

Journal entry for the above transaction

Tax Expense Dr. Deferred Tax Assets Dr. Tax Payable $147,000 $23,100 $170,100